OpenStudy (anonymous):

Write the slope-intercept equation for the line through (3, 7) and (9, 3).

hero (hero):

 Somebody likes making -2/3 the slope y - 3 = -2/3(x - 9) y = -2x/3 + 6 + 3 y = -2x/3 + 9